### v1

# Cold Beam of LiHe Dimers

Abstract
We demonstrate an approach for producing a high flux beam of LiHe dimers by performing post-nozzle seeding of a lithium beam into a supersonic helium expansion. The molecular beam has a transverse temperature of 11(4) mK and a continuous flux on the order of dimers/s. By simulating molecular formation in the jet, we find quantitative agreement with the measured dimer flux using ab-initio quantum scattering theory for three-body processes. Extensions of this work could lead towards a more quantitative understanding of other few-body collision processes of alkali and helium atoms.
Methods to produce cold molecules are of interest for studies of molecular collisions and ultra-cold chemistry, precision measurement experiments, and quantum science [1, 2]. In recent years, there has been great interest in three-body recombination as a fundamental chemical process [3, 4, 5, 6, 7]. A great deal of experimental work has been carried out [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], and quantitative theoretical understanding of these measurements is an important goal. Computing fully quantum ab-initio rate constants for complex molecules that contain a multitude of recombination channels remains a formidable challenge, but significant progress has been made in some systems [21, 19]. Simple molecules, particularly van der Waals (vdW) molecules containing hydrogen or helium, provide an alternative path towards quantitative understanding from first principles [22, 23, 24, 25, 26].
An intense cold atomic beam source that we have recently developed [27] can also be applied to quantitative studies of three-body processes. Using this source, we have produced a high flux beam of LiHe dimers with 11 mK temperature. We have also carried out Monte Carlo simulations of the molecular formation that incorporate currently available theoretical rate constants and found that it agrees with our measurements. Adopting our approach to other species opens up the possibility of further studies of van der Waals molecular formation and collisions in the 1 mK to 100 mK temperature range, intermediate between the sub-mK or above 100 mK temperatures of most previous work with cold molecules.
Various techniques have been used to produce cold molecules. Magneto-optical traps and laser cooling have been applied to a number of species and can achieve K temperatures [28, 16], but are not easily applicable to ground state helium . Buffer-gas based sources using helium have been widely used to create and cool a variety of molecules [29, 30, 15], including the first observation of LiHe by Tariq et al. [31]. Limited by the vapor pressure of helium, temperatures are typically in the single Kelvin range [29]. The CRESU method, based on seeded supersonic helium expansions, is also typically in the single to tens of Kelvin range [14].
In our approach, we make use of highly expanded supersonic helium, which can reach mK temperatures [32], enabling for example the first observation of He dimers and He trimers [33]. While the standard approach for producing seeded jets has been to perform seeding pre-expansion or at the opening of the jet nozzle [34, 35, 36, 37, 38], we instead perform post-expansion seeding [27]. This enables us to take full advantage of the high beam brightness of pure helium expansions while limiting heating on the helium jet by the seeded species as well as condensation on the nozzle.
Seeding the jet with lithium, a beam of LiHe is produced with the dimers forming primarily through three-body collisions Li + He + He LiHe + He. Laser-induced fluorescence on the single bound rovibrational ground state detects the dimers on the transitions, with . In contrast to previous work [31], the LiHe dimers are 100 times colder with a temperature of 11(4) mK.
The experimental apparatus for producing dimers is described in Ref [27]. To summarize, a supersonic helium expansion is produced by flowing helium gas through a nozzle cooled to about 4 K. The cryogenic temperature of the nozzle results in a modest forward velocity of the helium beam of about 210 m/s. Beneath the helium nozzle, an oven produces an effusive beam of lithium which is injected into the expanding jet. The lithium cools via collisions with the helium. A charcoal adsorption pump cooled to around 4 K surrounds the injection region and is used to remove background helium atoms. This allows for operation with continuous helium flow rates in excess of atoms/s along with continuous lithium flow rates into the jet of over atoms/s.
Fluorescence spectroscopy is used to characterize the cooling process of both the lithium and the dimers. This is performed with an external cavity diode laser which is locked to a low finesse etalon and has a mode-hop-free tuning range of about 6 GHz. An 871A Bristol wave-meter is used for frequency calibration of the molecular transitions. The wave-meter itself is calibrated against a reference fluorescence signal of the D1 and D2 lines of lithium produced in a separate beam source. The calibration and stability of the wave-meter currently limit the accuracy with which we can determine the absolute molecular transition frequencies.
An example fluorescence image along with spectral profiles of the transitions are shown in Fig. 1 for a distance of about 4 cm from the He nozzle. The signal to noise is , while the signal to noise for the transition, not shown, is . Absolute transition wavenumbers for both excited state transitions are given in Table 1. Due to the narrow velocity distribution in the moving frame of the jet, the hyper-fine transitions are resolvable. As the molecular beam expands spherically and as the probe laser is perpendicular to the center-line of the expansion, there are spatially dependent Doppler shifts. The hyper-fine transitions can thus be resolved as a function of frequency within a small region in space or at a single laser frequency as shown in Fig. 1 (a). Fitting the spectral profiles (black dots) in Fig. 1 (b) to Voigt profiles (red curves) allows us to determine the temperature of the dimers. At a distance of about 4 cm from the nozzle, the transverse temperature is 11(4) mK. In this analysis, it is assumed that the LiHe transitions have the same natural line-width of 5.87 MHz as the D1 line of Li [39].
From the fluorescence measurements, the peak density of the molecular beam at a given distance from the nozzle is calculated as,
where is the spatial profile of the molecular beam normalized to a height of one, is the laser intensity, is the resonant Doppler broadened cross-section for transitions between vibrational levels and , and is the number of scattered photons per second at resonance. The spatial profile is determined by fitting the fluorescence signal assuming cylindrical symmetry. The highest peak density occurs at the closest distance we can measure to the nozzle with a value of cm at a distance of cm.
| 14902.561 | 14881.615 | |
| 14902.563 | 14881.617 | |
| 14902.587 | 14881.642 | |
| 14902.589 | 14881.644 | |
| 14902.740 | 14881.778 | |
| 14902.768 | 14881.805 |
Evaluation of in Eq. 1 requires that we compute the Franck-Condon factors (FCF). We do so by numerically solving the radial Schrödinger equation for the ground and excited state potentials using the discrete variable representation (DVR) approach [40]. For the ground state potential, we use the potential from Ref. [25]. The excited state potential is treated as a Buckingham potential
where J m is taken from Ref. [41]. The coefficients and are determined by finding the values that reproduce the experimentally measured vibrational binding energies. In our fitting procedure, we neglect the fine, hyper-fine, and rotational structures as these splittings are small compared to the vibrational binding energies. The optimal values of the coefficients are found to be J and m.
All density measurements are acquired on the transition as it has the largest calculated FCF of . However, with this small of a FCF, most molecules will photo-dissociate upon absorption of a single photon. As a consequence, the laser intensity is kept sufficiently weak such that the probability of absorption of a photon as the molecules travel through the laser is .
The flux of dimers through a given area is computed as
where is the forward velocity of the molecular beam. For molecular results, we are unable to perform measurements of the longitudinal velocity distribution as scattered laser light off of the surface of the He nozzle saturates our camera. As such, we assume that the longitudinal velocity distribution of the dimers is similar to that of the lithium. That is a forward velocity in the lab frame of m/s while the root mean square velocity in the longitudinal direction is such that . The flux of dimers within a 2.54 cm diameter circle versus distance from the nozzle is shown in Fig. 2. Integrating over the entire spatial distribution of the beam yields a total flux of s. Based on measurements of the lithium flux entering the helium jet, about of the injected lithium atoms form into dimers. Studies at closer or further distances using the transverse laser probe are currently limited by optical access in the cryogenic region of our vacuum chamber.
A three-dimensional Monte-Carlo simulation is performed to compute the expected dimer flux. The simulation expands upon the method described in Ref. [43]. In the simulation, the velocity phase-space probability distribution of the helium jet is modeled from theory [44, 45], while the trajectories of the Li, LiHe, and formation and dissociation of the LiHe are treated with a Monte Carlo approach. Particles move in straight lines with time steps that are a small fraction of the local elastic collision rate, three-body recombination rate, or two-body loss rate. At each time step Li atoms and LiHe dimers may undergo two-body elastic collisions with He, Li atoms may undergo three-body inelastic collisions resulting in the formation of LiHe dimers, and LiHe dimers may undergo two-body inelastic collisions resulting in an unbound Li atom.
The rate equation governing the formation of dimers is
where and are the three-body formation rate and two-body loss rate respectively at a relative collision energy . Theoretical quantum calculations for for collision energies up to 10 mK have been performed by the authors of Ref. [23]. For large collision energies, we have examined two approaches. The first takes classical calculation from Ref. [42] which should be reliable above K. Modelling to decrease at a rate of between 10 mK and 1 K, smoothly connects these two regimes as shown in Fig. 3 (a). This approach does, however, likely underestimate at lower collision energies. In the second approach, we model to decrease at a rate of past 10 mK in accordance with the classical threshold law found in Ref. [46, 42]. For , we make use of theoretical calculations from Ref. [47] which include the processes LiHe + He Li + He + He and LiHe + He Li + He.
The two-body Li-He elastic collision rate is determined using calculated fully quantum differential cross-sections [25, 48]. For LiHe-He collisions, we make use of the elastic rate constant from Ref. [47] and assume isotropic scattering. We note that the theoretically calculated rate constants from Ref. [48, 47, 23, 42] use slightly different models for the ground state potential. At the moment, we are unable to quantify how these variations in the ground state potential affect the calculated rate constants and thus results from the simulations.
We treat scattering for three-body inelastic collisions and two-body collisions which remove LiHe as isotropic with momentum and energy being conserved. A binding energy of mK from Ref. [31] sets the amount of kinetic energy added or removed during the collisions. Shown in Fig. 2(a), this small change in kinetic energy amounts to a negligible change in particle’s trajectories post-collision. In fact, we find that varying the kinetic energy added by a factor of two results in just a few percent change in the simulation flux. As such, for computational simplicity, all loss collisions are treated as LiHe + He Li + He + He to determine the post-collision velocity of the Li atoms.
Simulated kinetic energy distributions of the Li and LiHe in the moving frame of the jet are presented in Fig. 3 (b). In contrast with the energy distribution of the LiHe, the Li distribution indicates a significant portion of Li atoms that have not fully thermalized with the helium. This behavior is observed experimentally in Ref. [27] with fluorescence measurements showing an asymmetric Doppler profile indicating a sizable fraction of lithium atoms that have not fully thermalized. Notably, the dimer fluorescence, as shown in Fig. 1 has no observable asymmetry. These results can be understood by examining the energy-dependent three-body recombination rate shown in Fig. 3 (a) alongside the energy distribution of the Li atoms. The portion of Li atoms with high kinetic energies simply have a drastically reduced probability of undergoing three-body collisions.
Results comparing simulated and experimentally measured dimer flux along with example particle trajectories are presented in Fig. 2. Four versions of the simulation are performed to explore how variations in rate constants affect the dimer flux. The “baseline” version of the simulation is the red triangles in Fig. 2 (b) with the other markers corresponding to changes in a single rate constant of the baseline version. For example, setting the loss rate to zero results in 7 times higher flux while removing LiHe-He collisions results in higher flux. The two approaches for modeling for mK, vary by . Agreement with experiment is highest when including LiHe-He collisions, dissociation collisions, and modeling to decrease at a rate of past 10 mK. For this model, the average fractional difference in flux between experiment and simulation is . When modeling to decrease as past 10 mK, the average fractional difference is 0.83(37). Overall, the experiment and model agree with each other to within about two standard deviations.
We note that our apparatus was not specifically designed to produce a high flux of dimers or to allow measurements over a very wide range of temperatures and densities. A modified apparatus could produce a higher flux of dimers and trimers over wider ranges of density and temperature. We hope to explore this in the future.
In summary, we have demonstrated an approach for producing a high flux beam of LiHe dimers with a temperature on the order of 10 mK. This method should be well suited for producing a number of ground state alkali-helium pairs given that theory predicts comparable recombination rates as Li + He + He LiHe + He [23]. Furthermore, all alkali-helium pairs have a large collisional cross-section with helium at low collision energies which is necessary for cooling and entrainment in the jet [22]. As such, our approach could facilitate quantitative studies of three-body recombination and other few-body collision processes for a variety of molecules including those yet to be observed experimentally.
We gratefully acknowledge Eite Tiesinga and Jacek Kłos, who shared their calculated Li-He cross sections with us. We acknowledge the financial support of this work by Fondren Foundation and the Army Research Laboratory Cooperative Research and Development Agreement number 16-080-004.