### v2

# Cold Beam of 7Li4He Dimers

Abstract
We demonstrate an approach for producing a high flux beam of 7Li4He dimers by performing post-nozzle seeding of a lithium beam into a supersonic helium expansion. The molecular beam has a longitudinal temperature of 13(6) mK and a continuous flux on the order of dimers/s. Monte Carlo simulations of molecular formation based on ab-initio quantum scattering calculations are carried out and compared to the experimentally observed molecular flux. Extensions of this work could lead towards a more quantitative understanding of few-body collision processes of alkali and helium atoms.
Methods to produce cold molecules are of interest for studies of molecular collisions and ultra-cold chemistry, precision measurement experiments, and quantum science [1, 2, 3]. In recent years, there has been great interest in molecules where fully quantum calculations are possible for studies of fundamental chemical processes such as three-body recombination [4, 5, 6, 7, 8]. Here, a great deal of experimental work has been carried out [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], and quantitative theoretical understanding of these measurements is an important goal.
Van der Waals (vdW) molecules containing hydrogen or helium are well suited for theoretical studies given their relatively simple structure [21, 22, 23, 24, 25]. For example, all alkali-helium pairs are predicted to contain a single bound-ground state [21] allowing for calculation of fully quantum ab-initio rate constants. More complex molecules containing a multitude of recombination channels makes such calculations challenging although progress has been made in some systems [26, 19]. Creating these alkali-helium pairs is not simple as the binding energies are predicted to be on the order of single to tens of mK [21, 24]. New methods to produce and study these molecules are in important step for connecting ab-initio calculations to experiment.
An intense cold atomic beam source that we have recently developed [27] can be used to produce alkali-helium dimers for studies of few-body processes. Using this source, we have produced a high flux beam of 7Li4He dimers with a longitudinal temperature of 13(6) mK. We have also carried out Monte Carlo simulations of the molecular formation that incorporate currently available theoretical rate constants. Adopting our approach to other species opens up the possibility of further studies of van der Waals molecular formation and collisions in the 1 mK to 100 mK temperature range, intermediate between the sub-mK or above 100 mK temperatures of most previous work with cold molecules.
Various techniques have been used to produce cold molecules. Magneto-optical traps and laser cooling have been applied to a number of species and can achieve K temperatures [28, 1], but are not easily applicable to ground state helium . Buffer-gas based sources using helium have been widely used to create and cool a variety of molecules [29, 30, 16], including the first observation of LiHe by Tariq et al. [31]. Limited by the vapor pressure of helium, temperatures are typically in the single Kelvin range [29]. The CRESU method, based on seeded supersonic helium expansions, is also typically in the single to tens of Kelvin range [15].
In our approach, we make use of highly expanded supersonic helium, which can reach mK temperatures [32], enabling for example the first observation of He2 dimers and He3 trimers [33]. While the standard approach for producing seeded jets has been to perform seeding pre-expansion or at the opening of the jet nozzle [34, 35, 36, 37, 38], we instead perform post-expansion seeding [27]. This enables us to take full advantage of the high beam brightness of pure helium expansions while limiting heating on the helium jet by the seeded species as well as condensation on the nozzle.
Seeding the jet with lithium, a beam of 7Li4He is produced as the 7Li becomes entrained and cooled in the helium jet. The main mechanisms which form the dimers are three-body recombination 7Li + 4He + 4He 7Li4He + 4He and two-body exchange collisions 7Li + 4He2 7Li4He + 4He. Laser-induced fluorescence on the single bound rovibrational ground state detects the dimers on the transitions, with . In contrast to previous work [31], the 7Li4He dimers are 100 times colder with a longitudinal temperature of 13(6) mK.
A diagram of the experimental apparatus is shown in Fig. 1 with further details described in Ref [27]. To summarize, a supersonic helium expansion is produced by flowing helium gas through a nozzle cooled to about 4 K. The cryogenic temperature of the nozzle results in a modest forward velocity of the helium beam of about m/s. Beneath the helium nozzle, an oven produces an effusive beam of lithium which is injected into the expanding jet. The lithium cools via collisions with the helium. A charcoal adsorption pump cooled to around 4 K surrounds the injection region and is used to remove background helium atoms. This allows for operation with continuous helium flow rates in excess of atoms/s along with continuous lithium flow rates into the jet of over atoms/s.
Fluorescence spectroscopy is used to characterize the cooling process of both the lithium and the dimers. This is performed with an external cavity diode laser which is locked to a low finesse etalon and has a mode-hop-free tuning range of about 6 GHz. An 871A Bristol wave-meter is used for frequency calibration of the molecular transitions. The wave-meter itself is calibrated against a reference fluorescence signal of the D1 and D2 lines of lithium produced in a separate beam source. The calibration and stability of the wave-meter currently limit the accuracy with which we can determine the absolute molecular transition frequencies.
| 14902.561 | 14881.615 | |
| 14902.563 | 14881.617 | |
| 14902.587 | 14881.642 | |
| 14902.589 | 14881.644 | |
| 14902.740 | 14881.778 | |
| 14902.768 | 14881.805 |
An example fluorescence image along with spectral profiles of the transitions are shown in Fig. 2 for a distance of about 4 cm from the 4He nozzle with a probe laser perpendicular to the expanding beam. Absolute transition wavenumbers for both excited state transitions are given in Table 1. Due to the narrow velocity distribution in the moving frame of the jet, the hyperfine transitions are resolvable. As the molecular beam expands spherically and as the probe laser is perpendicular to the center-line of the expansion, there are spatially dependent Doppler shifts. The hyperfine transitions can thus be resolved as a function of frequency within a small region in space or at a single laser frequency as shown in Fig. 2 (a). In Fig. 3 we show spectra over a smaller region in space of height 0.34 mm where geometric broadening is heavily reduced. Fitting the spectral profiles (black dots) in Fig. 3 to Voigt profiles (red curves) allows us to determine the local temperature of the dimers in the moving frame of the helium jet. At a distance of 4 cm from the nozzle, this temperature is 11(4) mK. In this analysis, it is assumed that the 7Li4He transitions have the same natural line-width of 5.87 MHz as the D1 line of 7Li [39]. We find that the separation of the excited state with hyperfine transtions between and is 70(5) MHz in contrast with the 7Li P1/2 hyperfine splitting of MHz [40]. The ground state hyperfine splitting between and is found to be 810(13) MHz, consistent with the 7Li ground state hyperfine splitting of MHz [40].
Longitudinal spectral profiles at a distance of about 3 cm from the nozzle are shown in Fig. 4. For longitudinal results, photo-dissociation effects as well as additional noise from scattered laser light reduces our ability to resolve signal at further distances. In addition it reduces the signal to noise compared to transverse spectra. However, at this distance, the longitudinal temperature is found to be 13(6) mK. Two additional peaks are shown in Fig. 4 compared to Fig. 2. This is due to reflections of the laser light off of the helium nozzle. The longitudinal probe beam produces spectral features which are blue detuned from resonance as the atoms are moving opposite to the laser propagation direction while the reflected probe beam produces spectral features which are red detuned. The additional peaks are the red detuned and transitions. Using the halfway point between the blue and red shifted Doppler features we estimate that the mean forward velocity of the dimer beam is 220(8) m/s.
From the transverse fluorescence measurements, the peak density of the molecular beam at a given distance from the nozzle is calculated as,
where is the spatial profile of the molecular beam normalized to a height of one, is the laser intensity, is the resonant Doppler broadened cross-section for transitions between vibrational levels and , and is the number of scattered photons per second at resonance. The spatial profile is determined by fitting the fluorescence signal assuming cylindrical symmetry. The highest peak density occurs at the closest distance we can measure to the nozzle with a value of cm-3 at a distance of cm.
Evaluation of in Eq. 1 requires that we compute the Franck-Condon factors (FCF). We do so by numerically solving the radial Schrödinger equation for the ground and excited state potentials using the discrete variable representation (DVR) approach [41]. For the ground state potential, we use the potential from Ref. [24]. The excited state potential is treated as a Buckingham potential
where J m6 is taken from Ref. [42]. The coefficients and are determined by finding the values that reproduce the experimentally measured vibrational binding energies. In our fitting procedure, we neglect the fine, hyperfine, and rotational structures as these splittings are small compared to the vibrational binding energies. The optimal values of the coefficients are found to be J and m.
All density measurements are acquired on the transition as it has the largest calculated FCF of . However, with this small of a FCF, most molecules will photo-dissociate upon absorption of a single photon. As a consequence, the laser intensity is kept sufficiently weak such that the probability of absorption of a photon as the molecules travel through the laser is .
The flux of dimers through a given area is computed as
where is the mean forward velocity of the molecular beam. The flux of dimers within a 2.54 cm diameter circle versus distance from the nozzle is shown in Fig. 5. Integrating over the entire spatial distribution of the beam yields a total flux of s-1. Based on measurements of the lithium flux entering the helium jet, about of the injected lithium atoms form into dimers. The molecules passing through a 2.54 cm diameter circle are confined to a solid angle
at 4 cm from the nozzle. Here is the full width at half maximum of the transverse velocity over the entire spatial distribution of the beam and is the mean forward velocity. Studies at closer or further distances than shown in Fig. 5 using the transverse laser probe are currently limited by optical access in the cryogenic region of our vacuum chamber.
A three-dimensional Monte-Carlo simulation is performed to compute the expected dimer flux. The simulation expands upon the method described in Ref. [44]. In the simulation, the velocity phase-space probability distribution of the helium jet is modeled from theory [45, 46], while the trajectories of the 7Li, 7Li4He, and formation and dissociation of the 7Li4He are treated with a Monte Carlo approach. Particles move in straight lines with time steps that are a small fraction of the local elastic collision rate, three-body recombination rate, two-body exchange rate, or two-body loss rate. At each time step 7Li atoms and 7Li4He dimers may undergo two-body elastic collisions with 4He, 7Li atoms may undergo three-body inelastic collisions or two-body exchange collisions resulting in the formation of 7Li4He dimers, and 7Li4He dimers may undergo two-body inelastic collisions resulting in an unbound 7Li atom.
The rate equation governing the formation of 7Li4He dimers is
where , and are the three-body formation rate, two-body exchange rate and two-body loss rate respectively at a relative collision energy . Theoretical quantum calculations for for collision energies up to 10 mK have been performed by the authors of Ref. [22]. For large collision energies, we have examined two approaches. The first takes classical calculation from Ref. [43] which should be reliable above K. Modelling to decrease at a rate of between 10 mK and 1 K, smoothly connects these two regimes as shown in Fig. 6 (a). This approach does, however, likely underestimate at lower collision energies. In the second approach, we model to decrease at a rate of past 10 mK in accordance with the classical threshold law found in Ref. [47, 43]. For , we make use of theoretical calculations from Ref. [48] which include the processes 7Li4He + 4He 7Li + 4He + 4He and 7Li4He + 4He 7Li + 4He2. The exchange rate for 7Li + 4He2 7Li4He + 4He is also taken from Ref. [48].
The two-body 7Li-4He elastic collision rate is determined using calculated fully quantum differential cross-sections [24, 49]. For 7Li4He-4He collisions, we make use of the elastic rate constant from Ref. [48] and assume isotropic scattering. We note that the theoretically calculated rate constants from Ref. [49, 48, 22, 43] use slightly different models for the ground state potential. At the moment, we are unable to quantify how these variations in the ground state potential affect the calculated rate constants and thus results from the simulations.
We treat scattering for three-body inelastic collisions and two-body collisions which remove 7Li4He as isotropic with momentum and energy being conserved. We assume the binding energy is mK based on the measured binding energy of mK from Ref. [31]. This sets the amount of kinetic energy added or removed during the collisions. Shown in Fig. 5(a), this small change in kinetic energy amounts to a negligible change in particle’s trajectories post-collision. In fact, we find that varying the kinetic energy added by a factor of two results in just a few percent change in the simulation flux. As such, for computational simplicity, all loss collisions are treated as 7Li4He + 4He 7Li + 4He + 4He to determine the post-collision velocity of the 7Li atoms.
Simulated kinetic energy distributions of the 7Li and 7Li4He in the moving frame of the jet are presented in Fig. 6 (b). In contrast with the energy distribution of the 7Li4He, the 7Li distribution indicates a significant portion of 7Li atoms that have not fully thermalized with the helium. This behavior is observed experimentally in Ref. [27] with fluorescence measurements showing an asymmetric Doppler profile indicating a sizable fraction of lithium atoms that have not fully thermalized. Notably, the dimer fluorescence, as shown in Fig. 2 has no observable asymmetry. These results can be understood by examining the energy-dependent three-body recombination rate shown in Fig. 6 (a) alongside the energy distribution of the 7Li atoms. The portion of 7Li atoms with high kinetic energies simply have a drastically reduced probability of undergoing three-body collisions.
As discussed below, we are currently unable to accurately determine the density of helium dimers in our system and their contribution to the formation of 7Li4He dimers. We have designed our source, as discussed in Ref. [27], to operate in a regime where helium cluster formation should be minimal. However, as exchange collisions are two-body process, this can be an important mechanism even with a 4He2 density that is a small fraction of the 4He density. We therefore explore a number of models to examine how variations in different rate constants affect the simulated dimer flux. Results from five versions of our simulations are shown Fig. 5. The “baseline” version of the simulation is the red triangles in Fig. 5 (b) in which . Other markers corresponding to changes in a single rate constant of the baseline version. For example, setting the loss rate to zero results in 7 times higher flux while removing 7Li4He-4He collisions results in higher flux. The two approaches for modeling for mK, vary by . Agreement with experiment is highest when including 7Li4He-4He collisions, dissociation collisions, and modeling to decrease at a rate of past 10 mK. For this model, the average fractional difference in flux between experiment and simulation is . When modeling to decrease as past 10 mK, the average fractional difference is 0.83(37).
Using a rate equation model discussed in the Appendix, we have estimated that the helium dimer density could be as high as of the helium density. This assumes a pure adiabatic expansion where heat of condensation contributes to negligible heating of the helium. Modelling the helium dimers as being in thermal equilibrium with the helium, we find that the inclusion of exchange collisions results in a 4-6 times increase in the 7Li4He flux over the range of distances that we measure. This would indicate that this is the dominant process over three-body formation. However, unlike the capture efficiency of 7Li and three-body formation of 7Li4He, the density of 4He2 dimers depends strongly on minor increases in the helium temperature on the order of a few milliKelvin. This is due to the helium binding energy being on the order of 1 mK, while the 7Li4He binding energy is on the order of 30 mK. Some results discussed in Ref. [27], show that minor temperature increased in the model of the helium temperature yield better agreement with experimental results of the seeded 7Li. Using a simple model in which a temperature floor is set in the helium expansion we repeat our simulations with a temperature floor of 1 mK and 5 mK. Results are shown in Fig. 7. A temperature floor of 1 mK reduces the 7Li4He flux by about 10 when including exchange collisions and only by a few percent (not shown) when . For a temperature floor of 5 mK, the flux including exchange collisions decreases by a factor of 3 to 4 while the flux with decreases by at most 20 . While this model does not capture the full dynamics of heating of the helium jet, it indicates that the contributions from exchange collisions have a strong temperature dependence.
Without a more physical model of our system that includes helium dimer formation, future studies will be necessary to determine quantitative rate constants. Monte-Carlo simulations of the entire helium expansion, helium dimer formation, and seeding of the lithium and formation of lithium-helium is one path forward but is prohibitive for our current computational resources. Experimentally, the helium dimer density could be determined with a transmission grating [11] while the helium temperature could be accessed using a pulsed electron beam to make meta-stable helium and performing time of flight measurements. We note that our apparatus was not specifically designed to produce a high flux of 7Li4He dimers or 4He2 dimers or to allow for measurements over a very wide range of temperatures and densities. A modified apparatus could produce a higher flux of dimers and trimers over a wider range of densities and temperatures and enable studies of three-body collisions as well as the possiblity of exchange collisions. We hope to explore this in future work.
In summary, we have demonstrated an approach for producing a high flux beam of 7Li4He dimers with a temperature on the order of 10 mK. This method should be well suited for producing a number of ground state alkali-helium pairs given that theory predicts comparable recombination rates as 7Li + 4He + 4He 7Li4He + 4He [22]. Furthermore, all alkali-helium pairs have a large collisional cross-section with helium at low collision energies which is necessary for cooling and entrainment in the jet [21]. For example, in Ref. [44], we simulate the seeding process for 87Rb and predict that of injected rubidium atoms could be entrained within a 0.02 sr solid angle with an energy of less than mK in the moving frame of the helium at a distance of 4 cm. This is roughly an order of magnitude less efficient than with lithium. Even with a dimer formation rate that is an order of magnitude less than for helium-lithium, this would roughly amount to a beam flux of dimers/s. As such, our approach could facilitate quantitative studies of three-body recombination and other few-body collision processes for a variety of molecules including those yet to be observed experimentally.
We gratefully acknowledge Eite Tiesinga and Jacek Kłos, who shared their calculated 7Li-4He cross sections with us. We also thank Paul Kunz for sharing lab equipment with us. We acknowledge the financial support of this work by Fondren Foundation and the Army Research Laboratory Cooperative Research and Development Agreement number 16-080-004.
I Appendix
We estimate the density of helium dimers as a function of distance from the nozzle along the centerline of the expansion by modeling the formation process as
Here and are the three-body recombination and two-body dissociation rates with brackets indicating a thermal average. The thermally averaged rate constants are computed using the energy-dependent recombination and dissociation rates from Ref. [50] and assuming that the helium and helium dimers are in thermal equilibrium and have a temperature which follows from Eq. 7.
In our model, we approximate the flow speed of the helium and helium dimers as being equal to the terminal velocity of the jet such that . This is due to the general feature that for supersonic expansions the flow speed approaches within a few percent of the terminal velocity within about five nozzle diameters [45]. For our source, the nozzle diameter is 200 microns. The final terms in the rate equations account for the density decreasing as a rate of with distance from the nozzle. For our nozzle geometry, the temperature of the helium after a distance of a few nozzle diameters is expected to decrease as
where is the reservoir temperature of the helium. Eq. 7 will hold in the limit that heat of condensation from dimer formation is negligible as well as a helium-helium collision rate which is sufficient to maintain local thermal equilibrium. For our source conditions, as discussed in Ref. [44], we expect the helium to remain highly collisional for distances exceeding 10 cm from our nozzle. Finally, for determining the dimer density off axis, the angular dependence of the helium density for moderate angles from the center-line of the expansion is treated as
Further discussions of the properties of supersonic expansions can be found in Ref. [45, 46].
Numerically solving Eq. 6 for our source conditions starting from four nozzle diameters yields a helium dimer density that is that of the helium past a few cm from the nozzle. The angular dependence at is shown in Fig. 8. Also shown in Fig. 8 is the simulated helium dimer density for a model of the helium expansion where a temperature floor of 5 mK is set. Compared to a pure adiabatic expansion, this temperature floor decreased the fraction of helium dimers by about a factor of 6 at 4 cm from the nozzle. While this is not a physically accurate way to model heat of condensation or heating from the lithium, it demonstrates the strong temperature dependence on the formation of helium dimers by minor increases to the helium temperature.